# Editable Fairness: Fine-Grained Bias Mitigation in Language Models

### v1

@@ Abstract
Generating fair and accurate predictions plays a pivotal role in deploying large language models (LLMs) in the real world. However, existing debiasing methods inevitably generate unfair or incorrect predictions as they are designed and evaluated to achieve parity across different social groups but leave aside individual commonsense facts, resulting in modified knowledge that elicits unreasonable or undesired predictions. In this paper, we first establish a new bias mitigation benchmark, BiaScope, which systematically assesses performance by leveraging newly constructed datasets and metrics on knowledge retention and generalization. Then, we propose a novel debiasing approach, Fairness Stamp (FAST), which enables fine-grained calibration of individual social biases. FAST identifies the decisive layer responsible for storing social biases and then calibrates its outputs by integrating a small modular network, considering both bias mitigation and knowledge-preserving demands. Comprehensive experiments demonstrate that FAST surpasses state-of-the-art baselines with superior debiasing performance while not compromising the overall model capability for knowledge retention and downstream predictions. This highlights the potential of fine-grained debiasing strategies to achieve fairness in LLMs. Code will be publicly available.
@@ 1 Introduction
Pre-trained Large Language Models (LLMs) have demonstrated exceptional performance on many tasks, such as language understanding and question answering [16, 21, 5]. However, the encoded social stereotypes and human-like biases inevitably cause undesired behaviors when deploying LLMs in practice [76, 54], e.g., making stereotyped judgments on vulnerable groups [62]. Removing such biases can not only enhance the generalization ability and reliability of LLMs but also expedite their deployment while retaining substantial social significance, which garners increasing attention from researchers, practitioners, and the broader public [45, 24, 36].
Existing approaches to mitigate biases in LLMs mainly fall into two categories: debiasing with fine-tuning or prompt-tuning [22, 23]. Fine-tuning methods include supplementary pre-training on a balanced corpus [78], aligning embeddings within the bias subspace [43, 58], or using a contrastive objective [33, 10]. Prompt-tuning techniques [27, 73, 42, 18] involve learning prompts that guide LLMs to ignore disparities between social groups for fair decision-making.
@@
However, these methods put emphasis on parity across different demographic groups, while generating unreasonable predictions on commonsense knowledge or are prone to exhibiting new biases regarding individual facts [22]. For instance, as shown in Figure 1, it is critical for debiased models to ensure fair correlations between social groups ‘male/female’ and descriptions ‘skilled at engineering/perform surgery’. However, as for individual facts such as “My [mom/dad] gives birth to me.”, it is not desirable if the model generates equitable predictions for social groups ‘mom/dad’, as the gender differences in these contexts are facts rather than stereotyped social biases [30, 22, 38, 15]. The problem could be two-fold. On one hand, existing debiasing approaches remove biases with group-invariant objectives [43, 33, 18], regarding different social groups as interchangeable. However, indiscriminately neutralizing different social groups degrades the perception of them, causing undesired or wrong behaviors when some individual statements hold distinct facts. On the other hand, existing datasets and benchmarks still lack sufficient evaluation regarding the effects of debiasing techniques. They mainly focus on evaluating the fairness on social biases, but ignore evaluating whether the debiased model retains commonsense knowledge or individual facts [22], for instance, reasonable disparities among different social groups.
@@
To address these issues, we first establish a new debiasing benchmark BiaScope, consisting of new datasets and metrics to further evaluate the effects of debiasing techniques. Specifically, BiaScope is established in two parts. To evaluate the ability to retain individual facts, we construct a dataset of commonsense knowledge about different social groups that should not be neutralized (e.g., My mom gives birth to me.). As a complement, we generate a dataset of paraphrased social biases to evaluate the generalization ability. In correspondence, we design two metrics, Retention Score (RS) and Paraphrase Stereotype Score (PS). Moreover, we propose a novel method Fairness-Stamp (FAST), for editable bias mitigation. Instead of mitigating group biases indiscriminately, FAST operates fine-grained calibrations on individual biases, i.e., specific biased description toward a social group [63, 12]. In particular, it operates through a two-stage pipeline, as illustrated in Figure 1: identifying where biases are encoded in LLMs and subsequently calibrating them by inserting a trainable fair stamp, i.e., a lightweight modular network, into the identified location. Aiming at bias mitigation and knowledge retention, we propose a novel weighted loss function to train the fair stamp while leaving all other model parameters frozen.
@@
We evaluate FAST with comprehensive experiments on StereoSet, Crows-Pairs, and our proposed BiaScope for systematic evaluation. Results show that FAST achieves superior debiasing performance without compromising model capability. Additional experiments showcase the scalability to larger models, the effectiveness on downstream tasks, and the effectiveness of knowledge locating, as well as analysis on fairness-utility trade-off and computational complexity. These underscore the immense potential of our fine-grained strategy in the realm of LLM debiasing. Our contributions are:
@@
  • We identify that existing debiasing benchmarks lack sufficient evaluation regarding the effects of debiasing techniques, and introduce a new benchmark BiaScope to evaluate the ability to retain individual commonsense facts and generalize to other social biases.
  • We propose a novel bias mitigation framework FAST, which calibrates fine-grained individual biases, enabling nuanced model debiasing.
@@
  • Comprehensive experiments and superior performance demonstrate the significant potential of our fine-grained debiasing strategy in achieving fairness for LLMs.
2 Related Works
Bias Mitigation in Large Language Models.
Several approaches have been proposed for debiasing pre-trained language models, which can be grouped into two categories: (1) Fine-tuning. This branch includes additional pre-training on re-balanced corpora [78, 70] or with a contrastive objective [33, 10], projection-based methods [43, 58, 36, 13] in the embedding space, in-training methods [29, 33] and parameter-efficient fine-tuning [39, 72] methods. (2) Prompt-tuning. Prompt-tuning [27, 73, 42, 18] involve generating either discrete prompts or continuous prompts to mitigate social biases. There are also post-hoc approaches [61, 7] that are deployed after the training phase to achieve effective debiasing. However, existing techniques treat social groups as interchangeable [22] and neutralize different social groups in model inputs or outputs, while ignoring or concealing distinct facts of different social groups [30]. In contrast, our method mitigates biases based on fine-grained individual biases, avoiding compromising other knowledge.
Knowledge Editing.
Knowledge or Model Editing [63, 12, 11] has been proposed to facilitate data-efficient modifications to model behavior while ensuring no detrimental impact on performance across other inputs. These approaches manipulate the model’s output for specific cases either by integrating external models with the original, unchanged model [49, 50, 17, 31, 34, 77], or by altering the model parameters responsible for undesirable output [48, 28, 32, 46]. The most relevant line of work is locate and edit [46, 47, 11, 41, 9], which suggests identifying neurons crucial to a model’s factual predictions [65, 20, 8] and subsequently updating the feed-forward weights to edit the output. Inspired by these works, we propose the first fine-grained bias mitigation framework, which enables nuancedly calibrating individual biases with minimal cost.
@@ 3 BiaScope Benchmark
Existing debiasing benchmarks focus on evaluating the fairness regarding social biases, while ignore evaluating the retention of commonsense knowledge [22]. In this paper, we establish the BiaScope benchmark, which includes new datasets and metrics designed for a more comprehensive evaluation of the modifications made by debiasing approaches. First, we describe the process of constructing datasets in Section 3.1. Then, we describe the corresponding evaluating metrics in Section 3.2.
@@ 3.1 Dataset Construction
The main idea of dataset construction is two-fold. First, to measure the ability of knowledge retention, we propose to create a commonsense knowledge dataset. Second, to prevent excessive knowledge retention and to measure generalization ability, we propose to construct a paraphrased social bias dataset. The process of dataset construction is illustrated in Figure 2. To ensure the quality of the generated data, we propose to collect real-world social biases $\Omega_{S}$ from existing datasets, which will serve as the basis for data generation. Social biases are gathered from three domains (gender, race, and religion) across six datasets, including StereoSet [51], Crows-Pairs [53], WEAT [6], WinoBias [75], Winogender [59], and BEC-Pro [2]. Each dataset comprises sentences or words demonstrating biases, with details provided in Appendix C.1.
@@
Create commonsense knowledge dataset. To better distinguish the boundary between out-of-scope knowledge and in-scope biases, we propose creating commonsense knowledge about sensitive groups. First, we extract sensitive subjects (e.g., man/woman, Christians/Jews) from $\Omega_{S}$. Then, we generate commonsense knowledge $\Omega_{R}$ about these subjects by prompting GPT-4. Finally, we manually validate the usability of $\Omega_{R}$. Knowledge in $\Omega_{R}$ does not constitute bias and should be retained after debiasing. However, it tends to be distorted by group-invariant debiasing methods.

### v2

# Identifying and Mitigating Social Bias Knowledge in Language Models

@@ Abstract
Generating fair and accurate predictions plays a pivotal role in deploying pre-trained language models (PLMs) in the real world. However, existing debiasing methods may inevitably generate incorrect or nonsensical predictions as they are designed and evaluated to achieve parity across different social groups but leave aside individual commonsense facts, resulting in modified knowledge that elicits unreasonable or undesired predictions. This paper introduces a novel debiasing framework that first identifies the encoding locations of biases within language models and then applies the Fairness-Stamp (FAST). FAST focuses on fine-grained, individual bias mitigation and integrates a lightweight network into PLMs, specifically targeting identified biases while preserving essential knowledge and maintaining factual integrity. We also present BiaScope, a new benchmark comprising datasets and metrics designed to evaluate the retention of commonsense knowledge and the generalization across paraphrased social biases. Our extensive experiments across multiple datasets demonstrate that FAST surpasses state-of-the-art baselines with superior debiasing performance while not compromising the overall model capability for knowledge retention and downstream predictions. This highlights the potential of fine-grained debiasing strategies to achieve fairness in PLMs. Code will be publicly available.
@@
Identifying and Mitigating Social Bias Knowledge in Language Models
Ruizhe Chen 1,5thanks: Equally Contributed.  Yichen Li 1,5  Jianfei Yang2 Yang Feng3  Joey Tianyi Zhou4 Jian Wu5Zuozhu Liu 1,5thanks: Corresponding author. 1ZJU-Angelalign R&D Center for Intelligence Healthcare, Zhejiang University 2Nanyang Technological University 3Angelalign Technology Inc. 4A*STAR Centre for Frontier AI Research 5Zhejiang University
@@ 1 Introduction
Pre-trained Language Models (PLMs) have demonstrated exceptional performance on many tasks, such as language understanding and question answering Devlin et al. (2018); Floridi and Chiriatti (2020); Brown et al. (2020). However, the encoded social stereotypes and human-like biases inevitably cause undesired behaviors when deploying PLMs in practice Zhao et al. (2019); Navigli et al. (2023), e.g., making stereotyped judgments on vulnerable groups Sheng et al. (2021). Removing such biases can not only enhance the generalization ability and reliability of PLMs but also expedite their deployment while retaining substantial social significance, which garners increasing attention from researchers, practitioners, and the broader public May et al. (2019); Gehman et al. (2020); Ma et al. (2023). Current approaches to mitigate biases in PLMs include debiasing through fine-tuning or prompt-tuning Gallegos et al. (2023); Garrido-Muñoz et al. (2021); Kaneko and Bollegala (2021). Fine-tuning involves additional pre-training on balanced corpora Zmigrod et al. (2019), aligning embeddings within bias subspaces Liang et al. (2020); Ravfogel et al. (2020), or using contrastive objectives He et al. (2022); Cheng et al. (2021) to lessen biases. Prompt-tuning techniques use prompts to guide PLMs towards ignoring social group disparities for fairer decision Guo et al. (2022); Yang et al. (2023); Li et al. (2023c); Dong et al. (2023).
@@
However, while these methods emphasize parity across different demographic groups, they also generate unreasonable predictions on commonsense knowledge and are prone to exhibiting new biases regarding individual facts Hanna et al. (2020); Gallegos et al. (2023); Kumar et al. (2022); Devinney et al. (2022). For example, as shown in Figure 1, for individual facts such as “The child is generally given birth by [mom/dad].”, applying parity indiscriminately incorrectly suggests both ‘mom’ and ‘dad’ could equally give birth, which biologically misrepresents factual differences and leads to nonsensical outcomes. This issue is caused by two factors. On one hand, existing debiasing approaches remove biases with group-invariant objectives Liang et al. (2020); He et al. (2022); Dong et al. (2023), regarding different social groups as interchangeable. However, individual statements hold distinct facts, while indiscriminately neutralizing different social groups degrades the perception of of individual facts, leading to undesired or wrong behaviors. On the other hand, current datasets and benchmarks primarily focus on assessing the fairness of social biases, but they do not adequately evaluate whether debiased models retain essential commonsense knowledge or respect factual differences among groups. These shortcomings may lead to models and methodologies that excessively prioritize equality at the cost of factual integrity Gallegos et al. (2023).
@@
To address these issues, we propose a novel framework, as illustrated in Figure 2. This framework focuses on identifying and mitigating individual social biases, rather than emphasizing group parity. In particular, we first formalize individual social bias as a knowledge, which is defined as a specific biased description toward a social group. Sinitsin et al. (2020); De Cao et al. (2021). Then, we identify where biases are encoded in language models by constructing counterfactual pairs with their unbiased alternatives. Finally, we introduce Fairness-Stamp (FAST), a novel approach that goes beyond indiscriminate mitigation of group biases. Unlike traditional methods, FAST performs fine-grained calibrations specifically targeting localized individual biases. FAST is designed as a learnable, lightweight modular network that is integrated into the identified location within the model. Its primary objectives are to mitigate biases while retaining other knowledge. Moreover, we establish a new debiasing benchmark, BiaScope, which includes newly created datasets and metrics designed to assess the effectiveness of various debiasing techniques in retaining factual knowledge. Specifically, BiaScope is established in two parts. First, to evaluate the ability to retain individual facts, we construct a dataset comprising commonsense knowledge about different social groups that should not be neutralized (e.g., My mom gives birth to me.). Second, to assess generalization capabilities, we have created a dataset of paraphrased social biases. Corresponding to these datasets, we have also designed two metrics: Retention Score (RS) and Paraphrase Stereotype Score (PS).
@@
We evaluate FAST with comprehensive experiments on StereoSet, Crows-Pairs, and our proposed BiaScope for systematic evaluation. The superior performance in bias mitigation and knowledge retention demonstrates the effectiveness of our framework in precisely identifying and calibrating social bias knowledge. Additional experiments showcase the scalability of larger models and the effectiveness of downstream tasks. Additional analysis showcases the effectiveness of knowledge localization, as well as analysis on fairness-utility trade-off and computational complexity. These underscore the immense potential of our fine-grained strategy in the realm of language model debiasing. Our contributions are:
  • Problem: We highlight an important problem where the excessive pursuit of equality between groups leads to incorrect predictions.
  • Algorithm: We propose a novel framework, FAST for this problem. Our framework identifies and mitigates fine-grained social bias knowledge.
@@
  • Dataset: We introduce a new benchmark, BiaScope, to evaluate the ability to retain individual commonsense facts and generalize to other social biases.
@@
  • Experiments: Our comprehensive experiments demonstrate superior performance, showcasing the effectiveness of our fine-grained debiasing strategy in enhancing fairness in language models.
Figure 2: An illustration of our framework: (a) We localize the bias knowledge that over-associates women with nurse than doctor in the language model. (b) We insert a fairness stamp to mitigate the bias knowledge at the localized layer.
2 Method
2.1 Preliminaries
Considering a transformer-based Language Model (specifically a decoder-only transformer), the model processes an input sequence $(x_{1},...,x_{t-1})$ and predicts the probability of the next token, denoted as $x_{t}$. The internal dynamics within a transformer block are expressed through the update of hidden states as follows:
$h_{t}^{(l)}$ $=h_{t}^{(l-1)}+\text{Attn}(h_{1}^{(l-1)},h_{2}^{(l-1)},\dots,h_{t}^{(l-1)})$
$\quad+\text{FFN}(h_{t}^{(l-1)}),$ (1)
where $h_{t}^{(l)}$ represents the hidden states at layer $l$, and the terms $\text{Attn}(\cdot)$ and $\text{FFN}(\cdot)$ signify the outputs from the self-attention layer and the feed-forward network layer at the $l^{th}$ level, respectively.
2.2 Social Bias Knowledge
Typically, a piece of social bias consists of a certain social group and the biased description that together amplify social inequalities Wang et al. (2023); Bommasani and Liang (2022); Allport et al. (1954). For instance, in the statement Mom is more likely to take care of the child., the phrase take care of the child is the biased description associated with the social group Mom. In light of this, we formalize the social bias as follows, inspired by Petroni et al. (2019); Jiang et al. (2020).
Definition 1.
A social bias can be formalized as a knowledge triplet $k=(s,r,o)$, where $s$ is the subject (i.e., Mom), $o$ is the object (i.e., take care of the child), and $r$ is the relation between them (i.e., is more likely to).
Based on the definition of social bias knowledge, we further explore the mechanisms by which language models exhibit social bias knowledge in their predictions. Inspired by previous studies Petroni et al. (2019), we make the following assumption:
Assumption 1.
Social bias knowledge can be stored implicitly in the parameters of a language model, as similarly as knowledge base.
Task Formulation.
In this section, we propose to identify and mitigate social bias knowledge in language models. The main idea is in two steps: (1) investigating if there are specific model parameters (i.e., hidden states) that play a more crucial role in storing social bias knowledge (Sec.2.3); (2) investigating how to mitigate the localized bias knowledge (Sec.2.4).
2.3 Social Bias Knowledge Localization
Contrastive Social Biases Localization.
To investigate how social bias $(s_{1},r_{1},o_{1})$ is stored as association between the social group and biased description, we propose to use its counterfactual knowledge $(s_{2},r_{2},o_{2})$ for contrast. This involves altering either the social group or biased description (e.g., changing Mom to Dad) to better probe these biased associations. Inspired by Meng et al. (2022a), our contrastive bias localization is performed in three runs:
(1) Biased run: We input the biased prompt $(s_{1},r_{1})$ into the model and collect all hidden states $\{h^{(l)}$ |||| $l\in[1,L]\}$, during a forward run towards biased prediction, where $L$ is number of layers.
(2) Counterfactual run: We input the counterfactual prompt $(s_{2},r_{2})$ to the model to modify the biased prediction. Hidden states will also change due to the alteration of the input subject.
(3) Restore biased states: To measure the effect of certain layer $\hat{l}$ on the biased prediction, we restore the biased states $h^{(\hat{l})}$ of $s_{1}$ and perform the forward run. Then we calculate the recovery degree of biased prediction, as detailed below.
Determine the decisive layer.
Denote the prediction probability on the object of the biased run as $P[o]$, and the probability of the counterfactual run as $P^{*}[o]$. In this way, the total biased effect (TE) can be defined as: TE $=P[o]-P^{*}[o]$. In the restoration run, the probability will recover from $P^{*}[o]$ to $P[o]$ due to the restoration of certain biased states $h^{(l)}$, which reflects the contribution of these states to the biased prediction. Denote the probability of restoring layer $l$ as $P^{*}(h^{(l)})[o]$. The indirect biased effect (IE), i.e., recovery degree, of layer $l$ can be calculated by IE $=P^{*}(h^{(l)})[o]-P^{*}[o]$. The layer demonstrating the largest IE is identified as the decisive layer.
2.4 Bias Mitigation with Fairness Stamp
Given a pre-trained language model $\mathcal{G}$ and a set of social biases $\Omega$ to be calibrated, the task involves producing an edited model $\mathcal{G}^{*}$ where social biases in $\Omega$ can be fairly predicted while other knowledge is retained as in the original $\mathcal{G}$. Following Sec. 2.3, we propose to envelop the decisive layer with an auxiliary fairness stamp (FAST), which can repair fine-grained social bias knowledge by editing a small number of weights.
Assuming the input hidden states to be $h$, the decisive layer (i.e., feed-forward network, FFN) in the original language model can be formulated as follows:
$\text{FFN}(h)=\text{Act}(h\mathbf{K}^{\top})\mathbf{V},$ (2)
where $\mathbf{K}$ and $\mathbf{V}$ denote the parameters (i.e., keys and values matrices) of the first and second linear layers in the FFN, respectively. The proposed fairness stamp is a 2-layer Feed-Forward Network (FFN), which helps modify the output of the decisive layer with a few external parameters to achieve the goal of fairness. The output of the enveloped FFN layer is given by:
$\text{FFN}^{\prime}(h)=\text{FFN}(h)+\text{Act}(h\mathbf{K^{\prime}}^{\top})\mathbf{V^{\prime}},$ (3)
Bias Mitigation.
With a social bias $k_{i}$ and its counterfactual knowledge $k_{i}^{{}^{\prime}}$, we propose to mitigate the gap between their probabilities of prediction on the associated objects:
$\mathcal{L}_{e}=\frac{1}{\lvert\Omega\rvert}\sum_{k_{i}\in\Omega}\lvert\mathcal{P}_{\mathcal{G}}[k_{i}]-\mathcal{P}_{\mathcal{G}}[k_{i}^{{}^{\prime}}]\rvert,$ (4)
where $k_{i}=(s_{i},r_{i},o_{i})$ follows the definition in Section 2.2. $\mathcal{P}_{\mathcal{G}}[k_{i}]=\mathcal{P}_{\mathcal{G}}[o_{i}|p_{i}]=\mathcal{P}_{\mathcal{G}}[o_{i}|s_{i},r_{i}]$ denotes the probability of predicting the object $o_{i}$ given the prompt $p_{i}$, where the prompt $p_{i}$ is composed of $s_{i}$ and $r_{i}$. Therefore, $k_{i}$ can also be expressed as $(p_{i},o_{i})$.
Knowledge Retention.
We aim to retain knowledge in two ways: firstly, by preserving the probability distribution of input prompts $p_{i}$ to minimize deviations from the original model. Second, we retain the probability distribution on the prompt $p^{\prime}$ that combines pre-defined template (e.g., “{subject} is _”) and the input subject (e.g., Mom), which helps retain the perception of different social groups and prevent the model from degradation of knowledge. The two loss functions are as follows:
$\mathcal{L}_{s1}=\frac{1}{\lvert\Omega\rvert}\sum_{(p_{i},o_{i})\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p_{i}],\mathcal{P}_{\mathcal{G}^{*}}[\star|p_{i}]),$
$\mathcal{L}_{s2}=\frac{1}{\lvert\Omega\rvert}\sum_{(s_{i},r_{i},o_{i})\in\Omega}\mathcal{D}_{KL}(\mathcal{P}_{\mathcal{G}}[\star|p^{\prime}],\mathcal{P}_{\mathcal{G}^{*}}[\star|p^{\prime}]),$
where $\mathcal{P}_{\mathcal{G}}[\star|p]$ is the predicted probability vector of all objects. $\mathcal{G}$ and $\mathcal{G}^{*}$ represent the origin and debiased model. $\mathcal{D}_{KL}(\cdot,\cdot)$ represents the Kullback-Leibler Divergence.
To prevent the model from overfitting to particular inputs, we utilize prefix texts $x_{j}$ to enhance generalization ability across various contexts Meng et al. (2022a). These prefix texts are randomly generated by the model, for instance, “My father told me that”, and are concatenated to the front of the prompts. The overall objective can be formulated as follows with hyperparameters $\alpha$ and $\beta$:
@@ 3 BiaScope Benchmark
Existing debiasing benchmarks focus on evaluating the fairness regarding social biases, while ignore evaluating the retention of commonsense knowledge Gallegos et al. (2023). In this paper, we establish the BiaScope benchmark, which includes new datasets and metrics designed for a more comprehensive evaluation of the modifications made by debiasing approaches. First, we describe the process of constructing datasets in Section 3.1. Then, we describe the corresponding evaluating metrics in Section 3.2.
@@ 3.1 Dataset Construction
The main idea of dataset construction is two-fold. First, to measure the ability of knowledge retention, we propose to create a commonsense knowledge dataset. Second, to prevent excessive knowledge retention and to measure generalization ability, we propose to construct a paraphrased social bias dataset. The process of dataset construction is illustrated in Figure 5. To ensure the quality of the generated data, we propose to collect real-world social biases $\Omega_{S}$ from existing datasets, which will serve as the basis for data generation. Social biases are gathered from three domains (gender, race, and religion) across six datasets. Each dataset comprises sentences or words demonstrating biases, with details provided in Appendix A.1.
Create commonsense knowledge dataset.
@@
To better distinguish the boundary between out-of-scope knowledge and in-scope biases, we propose creating commonsense knowledge about sensitive groups. First, we extract sensitive subjects (e.g., man/woman, Christians/Jews) from $\Omega_{S}$. Then, we generate commonsense knowledge $\Omega_{R}$ about these subjects by prompting GPT-4. Finally, we manually validate the usability of $\Omega_{R}$. Knowledge in $\Omega_{R}$ does not constitute bias and should be retained after debiasing. However, it tends to be distorted by group-invariant debiasing methods.
Create paraphrased social bias dataset.